For any data set of n values with standard deviation s, every value must be within of the mean. In a class of 19 students, the heights of the students have a mean of 66.6 inches and a standard deviation of 3.0 inches. The tallest student in class, a hopeful member of the basketball team, claims to be
78.6 inches tall. Could he be telling the truth?

Answer by rothauserc(4718)

P( X < 78.6 ) is determined by calculating the z-score
z-score = (78.6 - 66.6) / 3 = 4, therefore
P( X < 78.6 ) = 1.00 or 100%
now
P( X > or = 78.6 ) = 1.00 - P( X < 78.6 )
P( X > or = 78.6 ) = 1.00 - 1.00 = 0
The student is not telling the truth